OpenStudy (anonymous):

what is the domain and range of both these equations 1) 3x / 5-|7-4x| 2) square root of -1 / (3-x)(2-x)

OpenStudy (anonymous):

The domain is all of the values which can be put into the equation. For 1) can you see which values of x we would not be allowed to input?

OpenStudy (anonymous):

@Traxter my answer is 1. domain: [7/4, ~) range : (~ ,5] not so sure of it's right then the range of the second one, i don't know.

OpenStudy (anonymous):

For the first one, we can't have the denominator equal to zero. So solve 5-|7-4x|=0 to find which values of x will be excluded from the domain. See how you do with that.

OpenStudy (anonymous):

Domain= R-{1/2}

OpenStudy (anonymous):

since it is in p/q form denominator should not be equal to 0

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

I think you missed a value of x in the first question which would make the denominator 0. \[5-|7-4x|=0\] implies \[|7-4x|=5\] So \[7-4x=5\] and \[4x-7=5\] Hence the domain is R - {1/2,3}

OpenStudy (anonymous):

Oh....Yes...there are two cases...

OpenStudy (anonymous):

Indeed :) Now are you able to use the domain we have found to find the range?

OpenStudy (anonymous):

i forgot..that

OpenStudy (anonymous):

for range put y =3x / 5-|7-4x|

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

@Traxter ...wat abt Range..

OpenStudy (anonymous):

5y - 7y + 4xy = 3x -------------Case I

OpenStudy (anonymous):

Rearrange the equation so that you get x in terms of y (you will get 2 of them because of the modulus) and find the domain of y.

OpenStudy (anonymous):

-2y = 3x - 4xy -2y = x(3-4y) x = -2y/(3-4y) 3- 4y not equal 0 y not equal 3/4 -------- 1 st soln

OpenStudy (anonymous):

Case II 5y + 7y - 4xy = 3x 12y = x(3 + 4y) x = 12y/(3+4y) 3+ 4y not equal 0 y not equal -3/4

OpenStudy (anonymous):

@Traxter .... Do...u think mine is correct

OpenStudy (anonymous):

Yes :) Good work.

OpenStudy (anonymous):

so what's the range? the union of both of them ? @Traxter ?

OpenStudy (anonymous):

Yes, the range will be R - {the values of y we found we could not have}

OpenStudy (anonymous):

and how about for the second one ? @Traxter @Yahoo! is it going to be domain: (~,3)U(3,2)U(2,~) ? range: [0,~) ?

OpenStudy (anonymous):

For the second one we cannot have the denominator equal to zero, that is, we cannot have (3-x)(2-x)=0. So x cannot be 3 and x cannot be 2. Also, since we have the square root of negatives we will be in the complex plane. So the domain will be C-{2,3}

OpenStudy (anonymous):

yeah, and the range ?

OpenStudy (anonymous):

Have a go at working it out yourself, using the same method we used before. Find x in terms of y (or x in terms of f(x), whatever notation you're using) and see which values of y we can't have.

OpenStudy (anonymous):

i'm not sure. well i'm sure i did it wrong.

OpenStudy (anonymous):

@Traxter